# and Tight DP Auditing

### v2

@@ 5 Systematizing DP Auditing Research Based on their Operational Modes
In this section, we systematize the state-of-the-art work on DP auditing based on the modes of operation specific to DP auditing. Specifically, we extract the threat models used in the literature and introduce a new taxonomy to further break down these threat models into specific adversarial capabilities. Furthermore, we extract the types of privacy attacks and evaluation functions instantiated by auditing techniques. Finally, we discuss different techniques to calculate confidence levels for audits.
@@ 5.1 Threat Modeling
In the context of private machine learning, different threat models have been used to instantiate privacy attacks (e.g., black-box, white-box, etc.). For more fine-grained analysis, we break them down into specific capabilities – namely, model visibility and types of canary used by the attacks – which enables us to systematize research along these axes independently.
Model Visibility. When training ML models using DP-SGD, only the final trained model is released and visible to the adversary.111Except for specific settings, e.g., Federated Learning [76], where algorithms like DP-FedAvg and DP-FedSGD inherently disclose intermediate model updates. However, the theoretical privacy analysis of DP-SGD assumes that the intermediate models across all iterations are released as we do not know how to prove DP guarantees otherwise. Thus, early work assumes that the adversary has access to all intermediate models [84, 83]. More recently, the focus has shifted to auditing DP-SGD when only the final model output [92, 101, 7, 72] is available.
@@
Canaries. DP definitions assume a strong adversary that can in theory define the entire dataset given as input to a given mechanism (i.e., “dataset canary”). However, in the context of DP-SGD, auditing in this threat model typically entails training on so-called “pathological” datasets, which can destroy model utility [84] or may not reflect real-world threat models. While these considerations do not affect the theoretical analysis of DP-SGD, which provides guarantees against the worst-case adversary, they can inform model developers of the privacy properties satisfied by the specific model being built.
@@
To account for the use of canaries, in Algorithm 1, the Adversary defines the entire dataset (line 1) only in the dataset canary threat model. Whereas in the gradient and sample canary threat models, in line 1, the dataset is (possibly randomly) sampled from a “natural” distribution, e.g., CIFAR-10, instead. Additionally, in the gradient canary threat model, the DP-SGD algorithm is modified to allow the Adversary to provide gradient canaries associated with the input canaries through the auxiliary information, which will be inserted by the algorithm in place of those computed on the input canaries.
@@ 5.2 Attacks
When auditing DP implementations of simpler algorithms (e.g., Laplace mechanism or Report Noisy Max) [33, 17, 66, 87], the algorithm can directly analyze the outputs of the mechanism. However, when auditing more complex algorithms (e.g., DP-SGD), it becomes crucial to rely on privacy attacks to post-process the raw outputs from mechanisms into observations that can be analyzed more easily. We discuss these attacks and their role next.
@@
Membership Inference (MI). When auditing differentially private machine learning algorithms (e.g., DP-SGD, DP synthetic data), a common post-processing technique is membership inference (MI) [53, 84, 68, 51, 105, 116, 83, 92, 101, 71, 74, 20, 23, 111, 8, 30, 44, 7, 115, 6, 5, 85, 19, 112, 46] owing to its popularity as a privacy measurement tool for (non-private) ML models. In an MI attack, the adversary attempts to infer whether a given record was used to train the model. This strongly aligns with the add/remove adjacency where either a canary is added to the dataset or not.
@@
Attribute Inference (AI). Although MI is the most commonly used privacy attack for auditing DP-ML techniques, in specific circumstances, such as Label DP [47] and synthetic data [8], Attribute Inference (AI) has also been used [73, 51]. In AI, the adversary attempts to infer specific private attributes of a given record, given the record’s public attributes. To perform auditing, Malek et al. [73] infer the label associated with the canary sample in Label DP, while Houssiau et al. [51] infer sensitive attributes of a tabular canary given all other attributes for DP synthetic data. Overall, establishing a baseline attack success rate is typically more complicated for AI [51].
Reconstruction (Recon). Drawing on the theoretical bounds by Hayes et al. [50], reconstruction attacks (Recon) can also be used to audit DP implementations [72]. Here, the adversary attempts to generate canaries from the data domain given only the output of the mechanism and the canary distribution. Although this can be difficult for arbitrary canary distributions, [72] shows that the problem is tractable when the canaries are uniformly sampled from discrete distributions each containing canaries.
@@ 5.3 Evaluation Functions
Once simple scalar scores have been extracted from the mechanism outputs, there are different ways to fulfill the goal of the audit. While early work required the mechanism to be run multiple times (), to reduce computational overhead, recent work has focused on doing so in a single run (). In theory, the evaluation function can be specifically designed with the goal of the audit in mind—i.e., verification or estimation. However, in practice, evaluation functions are typically designed for the estimation goal, while verification is done by simply checking if the estimated parameters are consistent with the claimed DP guarantees – i.e., compute and check if and for verification. Next, we discuss estimation variants of different evaluation functions.
@@ 5.3.1 Multiple Runs
Distribution estimation. Alternatively, one can empirically estimate the distributions the observations are drawn from. For instance, Askin et al. [12] and Gorla et al. [49] use kernel density estimators and histogram estimators, respectively, to estimate the distributions of observations drawn from and . After estimating the distribution, they test the approximate DP condition directly as done in output-set optimization. Conversely, Kong et al. [61] compute the Hockey-stick222 [98]. or Rényi divergence depending on the claimed DP guarantees, while Huang et al. [52] estimate the variance of the distributions and estimate the privacy guarantees from the variance.
@@ 5.3.2 One Run
Mahloujifar et al. [72] and Xiang et al. [112] improve by auditing with, respectively, -DP and information theory. Specifically, Mahloujifar et al. [72] extend [101] to cover -DP, audit using Recon attacks, and improve the bound on attack success of Hayes et al. [50]. Xiang et al. [112] stick to MI but model the auditing task using information-theoretic principles (i.e., noisy channels).
Beyond fully connected neural networks and convolutional neural networks, recent work by Panda et al. [89] use the techniques from [101] to optimize canary design and audit large language models.
@@ 5.5 Remarks
In this section, we reviewed and extracted key details of auditing techniques, including the types of mechanisms and DP guarantees audited, privacy attacks and evaluation functions employed, and the progress made along the three desiderata discussed in Section 3.3 (i.e., efficiency, end-to-end-ness, and tightness).
While DP audits have used a wide range of techniques, very few techniques have achieved all three desired properties, and only for simple DP mechanisms. Next, we discuss the main research findings based on our systematic review of the DP auditing body of work.
6 Towards Efficient, End-to-End, and Tight DP Audits
We now discuss several key details of DP auditing that have been neglected by prior work, discover limiting factors to achieving all three desiderata, and identify novel directions for future work to focus on.
| Sub-Sampling Scheme | ||||||
| Poisson | Sampling WR | Sampling WOR | Shuffling | Balls-in-Bins | ||
| Adjacency | Add/Remove | [120] | [64] | |||
| Edit | [62, 64] | [99] | [120, 64] | |||
| Zero-Out | [27] | [27, 28] | [26] | |||
@@ 6.1 DP Guarantees & Adjacency
Naturally, the DP guarantee variant being audited is a fundamental requirement of any DP auditing algorithm, as per Definition 3.2. As evident from Table 1, while earlier work [33, 18, 53, 84, 51, 12, 67] focuses on simpler variants of DP such as -DP and ()-DP, more recent audits [83, 7, 72, 19, 112] do so on -DP and -GDP instead to achieve tight audits of mechanisms.
This is because simpler variants of DP may not always describe the privacy leakage of specific mechanisms optimally, i.e., the actual privacy leakage of the mechanism may be far from the theoretical privacy leakage expected from simple DP guarantees. In turn, audits are loose since DP auditing measures the actual privacy leakage from mechanisms. For instance, Nasr et al. [83] achieve tight audits of DP-SGD using the -DP guarantee, which they show models the actual privacy leakage of DP-SGD much more closely compared to ()-DP initially used for auditing [53, 84]. Therefore, recent audits of DP-SGD use the -DP or the related -GDP guarantee instead [83, 7, 72, 19, 112].
@@
Moreover, the DP guarantees and adjacency notions audited might not always be explicitly stated. This is the case for the adjacency notion in several key papers [44, 7, 52, 85, 11]. Furthermore, several results [83, 7, 72, 19, 112] are given with respect to ()-DP, although closer inspection of their methodology reveals that they are instead auditing the -DP or -GDP guarantees.
@@ 6.2 Sub-sampling Schemes
Another key detail when auditing mechanisms is the underlying sub-sampling scheme. From Table 1, we observe that the most common sub-sampling scheme audited is Poisson [53, 84, 83, 71, 92, 101, 8, 44, 72, 19, 112], owing to its use in the DP-SGD algorithm. However, recent work [4, 6] has also looked at auditing the shuffling scheme. Overall, while audits not involving sub-sampling schemes are known to be tight [33, 18, 12, 67, 83, 72, 112], tight audits involving sub-sampling schemes are much rarer [84, 8, 44], usually requiring additional assumptions, such as step-by-step auditing or simple models.
In this context, the main challenge is having tight DP guarantees of mechanisms employing sub-sampling schemes, i.e., knowing that the actual privacy leakage of mechanisms matches the theoretical privacy leakage expected from the DP guarantee. Next, we focus on the Gaussian Mechanism since sub-sampling schemes are most commonly used with that, e.g., in DP-SGD [1]. In Table 2, we summarize the sub-sampling schemes and adjacency notions for which tight DP guarantees are currently known. While tight guarantees are known for Poisson sub-sampling for any adjacency and Sampling WOR, Sampling WR, and Balls-in-Bins sampling for the Add/Remove, Edit, and Zero-out adjacencies, respectively, for the vast majority of sub-sampling scheme and adjacency pairs, tight guarantees remain unknown.
| Canary Type | ||||
| Sample | Gradient | Dataset | ||
| Model Visibility | Intermediate Models | [52] | [16, 84, 83, 101, 92, 74, 72, 112] | [84] |
| Final Model | [53, 84, 68, 116, 83, 101, 30, 44, 72, 85, 112, 89] | [19] | [84, 5, 11] | |
@@ 6.3 Threat Models
One elusive combination of desired properties for audits of modern complex mechanisms like DP-SGD is for them to be both end-to-end and tight. Specifically, Table 1 reveals that end-to-end audits of DP-SGD are not tight [53, 84, 83, 92, 101, 7, 72, 19] and viceversa [84, 83]. The only exception is the work by Feng et al. [44], who achieve both tight and end-to-end audits of DP-SGD, but only for simple logistic regression models.
To analyze this further, we break down end-to-end audits into specific adversarial capabilities (model visibility and canary type), as explained in Section 5.1. In Table 3, we highlight the threat models where state-of-the-art DP-SGD audits are tight. Generally, tight auditing in stronger models has been accomplished; however, while there has been significant work on the weakest (Final Model, Sample Canary) model, aka “Black-box,” less work has focused on the slightly stronger (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) models.333To ease presentation, we use the (Model Visibility, Canary Type) notation to quickly refer to the corresponding threat models.
@@ 6.5 Privacy Region & Confidence Level
The concept of privacy region has been used by auditors to both verify and estimate DP guarantees of the mechanisms under study. In Figure 2, we plot how the Gaussian mechanism () gets audited under both the ()-DP and -GDP variants. More precisely, in Figure 2(a), the privacy region is estimated from the mechanism and compared against the hypothetical guarantees of -DP, which the mechanism does not satisfy. Note the presence of large sections of the estimated privacy region that fall well outside the shaded theoretical privacy region for -DP. This suggests that the claimed DP guarantees are wrong and they can therefore be refuted experimentally.
@@ 7 Discussion & Conclusion
This paper systematized research on DP auditing using a systematic framework geared to categorize the underlying settings/techniques and measure progress in the state of the art along three key desired properties: efficiency, end-to-end-ness, and tightness.
Overall, our analysis showed that while DP auditing has made significant and rapid progress in making audits more effective, useful, and easily deployable against production systems, there are still areas with substantial research gaps. In the process, we identified several key insights and open problems; we are confident that future work can build upon these to establish a “Gold Standard” for DP auditing.
In the rest of this section, we review related work, present actionable recommendations, and highlight open research questions future work should prioritize.
7.1 Related Work
To the best of our knowledge, our work is the first to systematically review and analyze the body of work on DP auditing. Namatevs et al. [81] recently present a survey of DP auditing research; however, they do not analyze relevant work within a systematic framework. By contrast, we do so to comprehensively analyze the state of the art, measure progress, identify gaps, and highlight open research problems primarily along three key characteristics: whether audits are efficient, end-to-end, and tight.
Systematization of knowledge efforts in related areas include [91], who present attacks and defenses on (non-private) machine learning systems using a comprehensive adversarial framework. Salem et al. [97] focus on privacy risks and provide a unifying framework relying on privacy games. Finally, Desfontaines and Pejó [32] systematize the various variants and extensions of DP. While DP auditing relies on and is informed by the notions of privacy attacks, games, and DP variants, our work focuses on the specific process of deriving empirical bounds, which has thus far not been explored in prior work.
@@ 7.2 What is considered SOTA in DP Auditing?
Although there are different settings and thus different auditing techniques in literature, we attempt to identify the results that can be considered state-of-the-art (SOTA) with respect to tightness. More precisely, in Table 4, we do so while breaking down the DP guarantees audited and whether the other two key desirable properties are achieved.
When audits can be inefficient but minimal utility loss is expected, we consider the auditing method by Jagielski et al. [53] to be SOTA for complex algorithms like DP-SGD and Lokna et al. [67] for simple mechanisms satisfying ()-DP. We also note that Chadha et al. [20]’s method is the only technique available for mechanisms satisfying -RDP. Finally, for -DP (or -GDP), Annamalai et al. [7]’s auditing method is SOTA for end-to-end audits, while that by Cebere et al. [19] is SOTA for audits over sub-sampling that do not need to be end-to-end.

### v3

@@ 5 Systematizing DP Auditing Research Based on their Operational Modes
In this section, we systematize the state of the art on DP auditing based on the modes of operation specific to it. Specifically, we extract the threat models used in the literature and introduce a new taxonomy to further break down these threat models into specific adversarial capabilities. Furthermore, we extract the types of privacy attacks and evaluation functions instantiated by auditing techniques. Finally, we discuss different techniques to calculate confidence levels for audits.
@@ 5.1 Threat Modeling
In the context of private machine learning, different threat models have been used to instantiate privacy attacks (e.g., black-box, white-box, etc.). For more fine-grained analysis, we break them down into specific capabilities, i.e., model visibility and types of canary used by the attacks. This enables us to systematize research along these axes independently.
Model Visibility. When training ML models using DP-SGD, only the final trained model is released and visible to the adversary.333Except for settings like Federated Learning [79], where algorithms like DP-FedAvg and DP-FedSGD inherently disclose intermediate model updates. However, the theoretical privacy analysis of DP-SGD assumes that the intermediate models across all iterations are released as we do not know how to prove DP guarantees otherwise. Thus, early work assumes that the adversary has access to all intermediate models [88, 87]. More recently, the focus has shifted to auditing DP-SGD when only the final model output [96, 105, 8, 75] is available.
@@
Canaries. DP definitions assume a strong adversary that can in theory define the entire dataset given as input to a given mechanism (i.e., “dataset canary”). However, in the context of DP-SGD, auditing in this threat model typically entails training on so-called “pathological” datasets, which can destroy model utility [88] or may not reflect real-world threat models. While these considerations do not affect the theoretical analysis of DP-SGD, which provides guarantees against the worst-case adversary, they can inform of the privacy properties satisfied by the specific model being built.
@@
To account for the use of canaries, in Algorithm 1, the Adversary defines the entire dataset (line 1) only in the dataset canary threat model. Whereas in the gradient and sample canary threat models, in line 1, the dataset is (possibly randomly) sampled from a “natural” distribution, e.g., CIFAR-10, instead. Additionally, in the gradient canary model, the DP-SGD algorithm is modified to allow the Adversary to provide gradient canaries associated with the input canaries through the auxiliary information, which will be inserted by the algorithm in place of those computed on the input canaries.
@@ 5.2 Attacks
When auditing DP implementations of simpler algorithms with low-dimensional output spaces (e.g., Laplace mechanism or Report Noisy Max) [34, 18, 69, 90], the DP auditing algorithm can directly analyze the outputs of the mechanism. However, auditing more complex mechanisms (e.g., DP-SGD) with high-dimensional output spaces can be especially challenging, as it is unclear how to transform the high-dimensional output to a privacy parameter estimate. One popular way is to rely on privacy attacks to post-process the raw outputs from mechanisms into observations that can be thresholded and analyzed more easily. We discuss these attacks and their role next.
@@
Membership Inference (MI). When auditing differentially private machine learning algorithms (e.g., DP-SGD, differentially private synthetic data), a common post-processing technique is membership inference (MI) [55, 88, 71, 53, 109, 120, 87, 96, 105, 74, 77, 22, 25, 115, 9, 32, 46, 8, 119, 7, 6, 89, 21, 116, 48] owing to its popularity as a privacy measurement tool for (non-private) ML models. In an MI attack, the adversary attempts to infer whether a given record was used to train the model. This strongly aligns with the add/remove adjacency where either a canary is added to the dataset or not.
@@
Attribute Inference (AI). Although MI is the most common privacy attack for auditing DP-ML techniques, in specific circumstances, such as Label DP [49] and synthetic data [9], Attribute Inference (AI) has also been used [76, 53]. In AI, the adversary attempts to infer specific private attributes of a given record, given the record’s public attributes. To perform auditing, Malek et al. [76] infer the label associated with the canary sample in Label DP, while Houssiau et al. [53] infer sensitive attributes of a tabular canary given all other attributes for DP synthetic data. Overall, establishing a baseline attack success rate is typically more complicated for AI [53].
Reconstruction (Recon). Drawing on the theoretical bounds by Hayes et al. [52], reconstruction attacks (Recon) can also be used to audit DP implementations [75]. Here, the adversary attempts to generate canaries from the data domain given only the output of the mechanism and the canary distribution. Although this can be difficult for arbitrary canary distributions, Mahloujifar et al. [75] show that the problem is tractable when the canaries are uniformly sampled from discrete distributions each containing canaries.
@@ 5.3 Evaluation Functions
Once simple scalar scores have been extracted from the mechanism outputs, there are several ways to achieve the audit’s goal. While early work requires running the mechanism multiple times (), to reduce computational overhead, recent work has focused on doing so in a single run (). In theory, the evaluation function can be specifically designed with the audit’s goal in mind, i.e., estimation or verification. In practice, however, evaluation functions are typically designed for the former, while the latter is done by simply checking if the estimated parameters are consistent with the claimed DP guarantees – i.e., compute and check if and for verification. Next, we discuss estimation variants of different evaluation functions.
@@ 5.3.1 Multiple Runs
Distribution estimation. Alternatively, one can empirically estimate the distributions the observations are drawn from. For instance, Askin et al. [13] and Gorla et al. [51] use kernel density estimators and histogram estimators, respectively, to estimate the distributions of observations drawn from and . After estimating, they test the approximate DP condition directly as done in output-set optimization. Conversely, Kong et al. [63] compute the Hockey-stick444 [102]. or Rényi divergence depending on the claimed DP guarantees, while Huang et al. [54] estimate the variance of the distributions and estimate the privacy guarantees from the variance.
@@ 5.3.2 One Run
Mahloujifar et al. [75] and Xiang et al. [116] improve by auditing with, respectively, -DP and information theory. Specifically, Mahloujifar et al. [75] extend [105] to cover -DP, audit using Recon attacks, and improve the bound on attack success of Hayes et al. [52]. Xiang et al. [116] stick to MI but model the auditing task using information-theoretic principles (i.e., noisy channels). Beyond fully connected neural networks and convolutional neural networks, recent work by Panda et al. [93] uses the techniques from [105] to optimize canary design and audit large language models.
@@ 5.5 Remarks
Overall, while DP audits have used a wide range of techniques, very few audits have achieved all three desired properties, and only for simple DP mechanisms.
6 Efficient, End-to-End, and Tight DP Audits?
Based on our systematic review of the DP auditing body of work, we now discuss several key details of DP auditing that have been neglected by prior work, discover limiting factors to achieving all three desiderata, and identify novel directions for future work to focus on.
@@ 6.1 DP Guarantees & Adjacency
Naturally, the DP guarantee variant being audited is a fundamental requirement of any DP auditing algorithm, as per Definition 3.2. As evident from Table 1, while earlier work [34, 19, 55, 88, 53, 13, 70] focuses on simpler variants like -DP and ()-DP, more recent audits [87, 8, 75, 21, 116] do so on -DP and -GDP to achieve tight audits.
This is because simpler variants of DP (-DP and ()-DP) may not always describe the privacy leakage of specific mechanisms optimally, i.e., the actual privacy leakage of the mechanism may be far from the theoretical privacy leakage expected from simple DP guarantees. In turn, audits are loose since DP auditing measures the actual privacy leakage from mechanisms. For instance, Nasr et al. [87] achieve tight audits of DP-SGD using the -DP guarantee, which they show models the actual privacy leakage of DP-SGD much more closely compared to ()-DP initially used for auditing [55, 88]. Therefore, recent audits of DP-SGD use the -DP or the related -GDP guarantee instead [87, 8, 75, 21, 116].
@@
Moreover, the DP guarantees and adjacency notions audited might not always be explicitly stated or are left unclear. This is the case for the adjacency notion in several key papers [46, 8, 54, 89, 12]. Furthermore, several results [87, 8, 75, 21, 116] are given with respect to ()-DP, although closer inspection of their methodology reveals that they are instead auditing the -DP or -GDP guarantees.
@@ 6.2 Sub-sampling Schemes
Another key detail in DP auditing is the underlying sub-sampling scheme. From Table 1, we observe that the most common sub-sampling scheme audited is Poisson [55, 88, 87, 74, 96, 105, 9, 46, 75, 21, 116], owing to its use in the DP-SGD algorithm. However, recent work [5, 7] has also looked at auditing the shuffling scheme. Overall, while audits not involving sub-sampling schemes are known to be tight [34, 19, 13, 70, 87, 75, 116], tight audits involving sub-sampling schemes are much rarer [88, 9, 46], usually requiring additional assumptions, such as step-by-step auditing or simple models.
In this context, the main challenge is having tight DP guarantees of mechanisms employing sub-sampling schemes, i.e., knowing that the actual privacy leakage of mechanisms matches the theoretical privacy leakage expected from the DP guarantee. Next, we focus on the Gaussian Mechanism since sub-sampling schemes are most commonly used in conjunction with that, e.g., in DP-SGD [1]. In Table 2, we summarize the sub-sampling schemes and adjacency notions for which tight DP guarantees are currently known. While tight guarantees are known for Poisson sub-sampling for any adjacency and Sampling WOR, Sampling WR, and Balls-in-Bins sampling for the Add/Remove, Edit, and Zero-out adjacencies, respectively, for the vast majority of sub-sampling scheme and adjacency pairs, tight guarantees remain unknown.
| Poisson | Sampling | Sampling | Shuffling | Balls-in- | |
| WR | WOR | Bins | |||
| Add/Remove | [125] | [67] | |||
| Edit | [64, 67] | [103] | [125, 67] | ||
| Zero-Out | [29] | [29, 30] | [28] |
@@ 6.3 Threat Models
One elusive combination of desired properties for audits of complex mechanisms like DP-SGD is for them to be both end-to-end and tight. Specifically, Table 1 reveals that end-to-end audits of DP-SGD are not tight [55, 88, 87, 96, 105, 8, 75, 21] and viceversa [88, 87]. The only exception is the work by Feng et al. [46], who achieve both tight and end-to-end audits of DP-SGD, but only for simple logistic regression models.
| Sample | Gradient | Dataset | |
| Intermediate | [54] | [17, 88, 87, 105, 96, 77, 75, 116] | [88] |
| Final | [55, 88, 71, 120, 87, 105, 32, 46, 75, 89, 116, 93] | [21] | [88, 6, 12] |
To analyze this further, we break down end-to-end audits into specific adversarial capabilities (model visibility and canary type) as explained in Section 5.1. In Table 3, we highlight the threat models where state-of-the-art DP-SGD audits are tight. Generally, tight auditing in stronger threat models has been accomplished; however, while there has been significant work on the weakest (Final Model, Sample Canary) (aka “Black-box”) threat model, less work has focused on the slightly stronger (Final Model, Gradient Canary) and (Intermediate Models, Sample Canary) models.555To ease presentation, we use the (Model Visibility, Canary Type) notation to quickly refer to the corresponding threat models.
@@ 6.5 Privacy Region & Confidence Level
The concept of privacy region has been used by auditors to both verify and estimate DP guarantees of the mechanisms under study. Figure 2 visualizes how the Gaussian mechanism () gets audited under both the ()-DP and -GDP variants. More precisely, in Figure 2(a), the privacy region is estimated from the mechanism and compared against the hypothetical guarantees of -DP, which the mechanism does not satisfy. Note the presence of large sections of the estimated privacy region that fall well outside the shaded theoretical privacy region for -DP. This suggests that the claimed DP guarantees are wrong and they can therefore be refuted experimentally.
@@ 7 Discussion & Conclusion
This paper systematized research on DP auditing using a systematic framework geared to categorize the underlying settings/techniques and measure progress in the state of the art along three key desired properties: efficiency, end-to-end-ness, and tightness. Overall, our analysis showed that while DP auditing made significant and rapid progress in making audits more effective, useful, and easily deployable on production systems, there are still areas with substantial research gaps. In the process, we identified several key insights and open problems; we are confident that future work can build upon these to establish a “Gold Standard” for DP auditing.
In the rest of this section, we present actionable recommendations, and highlight open research questions to prioritize.
@@ 7.1 What is considered SOTA in DP Auditing?
Although there are different settings and thus different auditing techniques in the literature, we attempt to identify the results that can be considered state-of-the-art (SOTA) with respect to tightness. To do so, we compare the reported empirical values at standardized theoretical levels (i.e., ) and algorithms (i.e., training CNN on CIFAR-10 [55, 88, 87, 105, 8, 21, 75, 116] when auditing DP-SGD). More precisely, in Table 4, we do so while breaking down the DP guarantees audited and whether the other two key desirable properties are achieved.
When audits can be inefficient but minimal utility loss is expected, we consider the method by Jagielski et al. [55] to be SOTA for complex algorithms like DP-SGD and Lokna et al. [70] for simple mechanisms satisfying ()-DP. Also, Chadha et al. [22]’s method is the only technique available for mechanisms satisfying -RDP. Finally, for -DP (or -GDP), Annamalai et al. [8]’s auditing method is SOTA for end-to-end audits, while that by Cebere et al. [21] is SOTA for audits over sub-sampling that do not need to be end-to-end.